Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

its an easy question

Expert replies
by sana.noor » Mon Aug 05, 2013 10:44 am
Tony’s political science final exam consists of eight True/False questions. If Tony guesses on every question, what is the probability that he gets exactly seven questions right?

A) 1/32
B) 1/16
C) 1/8
D) 7/8
E) 31/32

OA is A
Work hard in Silence, Let Success make the noise.

If you found my Post really helpful, then don't forget to click the Thank/follow me button. :)
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Mon Aug 05, 2013 11:53 am
sana.noor wrote:Tony’s political science final exam consists of eight True/False questions. If Tony guesses on every question, what is the probability that he gets exactly seven questions right?

A) 1/32
B) 1/16
C) 1/8
D) 7/8
E) 31/32

OA is A
P(exactly 7 right) = P(exactly 1 wrong).

P(exactly N times) = P(one way) * total possible ways.

One way:
One way to get exactly 1 wrong is to answer only the 1st question wrong.
P(WRRRRRRR) = 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 = 1/2�.

Total possible ways:
WRRRRRRR represents ONE WAY to answer exactly 1 question wrong.
To account for ALL OF THE WAYS to answer exactly 1 question wrong, the result above must be multiplied by the number of ways to arrange the letters WRRRRRRR.
Since W could appear in any of the 8 positions of the arrangement, the total number of ways to answer exactly 1 question wrong = 8.

Thus;
P(exactly 1 wrong) = (1/2�) * 8 = 1/32.

The correct answer is A.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by [email protected] » Mon Aug 05, 2013 11:55 am
Hi sana.noor,

Since this is a probability question, we'll use the probability formula:

(Desired Outcomes)/(Total Outcomes)

The test has 8 True/False questions and Tony is guessing on each, so there are 2^8 possible outcomes = 256 possible outcomes

For Tony to get exactly 7 out of 8 correct, it can be ANY combo of 7, so we'll use the combination formula:

8c7 = 8!/[(7!)(1!) = 8 combos of seven correct answers

8/256 = 1/32


GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion